OpenStudy (anonymous):

The school taxes on a house assessed at $210,000 are $2340. At the same tax rate, what are the taxes (to the nearest dollar) on a house assessed at $275,000? a. $3,064 b. $65,000 c. $1,787 d. $90

OpenStudy (tkhunny):

Have you considered a Proportion?

OpenStudy (anonymous):

2340/210000 =x/275000. Tax/home value = tax/home value

OpenStudy (anonymous):

2340 (275000) = 210000x. Solve for x

OpenStudy (tkhunny):

Now, see, this is why I don't like this - whatever you want to call it. You had this: 1) \(\dfrac{2340}{210000} = \dfrac{x}{275000}\) You did whatever it is you want to call it to obtain: 2) \(2340*275000 = x*210000\) And the last item is to solve for x, giving: 3) \(\dfrac{2340*275000}{210000} = x\) My question is this, we started with 210000 in the denominator (1). Why in the name of reason did we bother to put that 210000 up in the numerator of the other side (2)? We WANT it in that denominator (3)! Move it just to put it back? Why does this make sense to anyone?! This proportion business needs some reworking in our textbooks! My opinion.

OpenStudy (anonymous):

any answer that yall recommend

OpenStudy (tkhunny):

Yes. We both recommended "solve for x". Please do that and tell us your results.

OpenStudy (anonymous):

how do u solve for x

OpenStudy (tkhunny):

Look at my equation #3. If you cannot complete the problem from there, you need to back up a couple of chapters and pay better attention.